OpenStudy (anonymous):

Graph f(x)=x^2-3x-10/x-2. Find the domain, range, VA, HA/SA

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hey:)

OpenStudy (misty1212):

is it \[\frac{x^2-3x-10}{x-2}\]

OpenStudy (anonymous):

Yes

OpenStudy (misty1212):

ok lets do them one at a time

OpenStudy (misty1212):

domain is easy enough, set the denominator equal to zero and solve then say "all real numbers except that one"

OpenStudy (misty1212):

what do you get for that? (you can do it in your head)

OpenStudy (anonymous):

2

OpenStudy (misty1212):

yeah so all numbers except 2

OpenStudy (misty1212):

now before we continue, lets make sure this is not a trick question can you factor the numerator and cancel?

OpenStudy (anonymous):

Let me see...

OpenStudy (misty1212):

hmmi think it does factor, but not cancel !

OpenStudy (anonymous):

(x-5)(x+2)/x-2

OpenStudy (misty1212):

yeah but nothing cancels, so lets continue but you should always check that first

OpenStudy (misty1212):

once we have the domain, we also have the vertical asymptote too

OpenStudy (misty1212):

clear or no?

OpenStudy (anonymous):

So it is also 2

OpenStudy (anonymous):

yes!

OpenStudy (misty1212):

careful here

OpenStudy (misty1212):

the vertical asymptote is a vertical line, not a number

OpenStudy (misty1212):

so you want to return and answer of \(x=2\) the vertical line, not just the number 2

OpenStudy (misty1212):

clear?

OpenStudy (anonymous):

Oh! Yes! Clear

OpenStudy (misty1212):

now for the horizontal asympote the degree of the numerator is ?

OpenStudy (anonymous):

2

OpenStudy (misty1212):

and the denominator?

OpenStudy (anonymous):

there is no horizontal then since it's greater than the denom.

OpenStudy (misty1212):

bingo

OpenStudy (misty1212):

but the degree of the numerator is one more than the degree of the denominator that means there will be a slant asymptote do you know how to find it?

OpenStudy (anonymous):

I never learned about a slant asymptote so I dont know :(

OpenStudy (misty1212):

ok it is easy to say what to do , but not easy for me to write in one word "divide"

OpenStudy (misty1212):

when you divide a polynomial of degree 2 by a polynomial of degree 1 you get a polynomial of degree 1 ( a line) and a remainder ignore the remainder, the line is the slant asymptote

OpenStudy (misty1212):

you can use either long division or (much easier) synthetic division do you know how to do that?

OpenStudy (anonymous):

Yes :)

OpenStudy (misty1212):

whew like i said it is very hard for me to write division here what do you get?

OpenStudy (misty1212):

take your time

OpenStudy (anonymous):

I did it in a rush so I got x^2-1x-6.

OpenStudy (misty1212):

hmm no you should have a polynomial of degree 1, not a quadratic

OpenStudy (misty1212):

1 -3 -10 2 ______________ 1

OpenStudy (misty1212):

that is the iniital set up then 1 -3 -10 2 2 ______________ 1 -1

OpenStudy (misty1212):

the next line is unimportant since we only need the line, not the remainder the line has the coefficients from the final row so \(x-1\) slant asymptote is \(y=x-1\)

OpenStudy (anonymous):

Oh! Now I know what I did wrong! When did the fak divison, i fogot to add the "0^2" Lol but yes, I know understand

OpenStudy (misty1212):

ok then we are done!

OpenStudy (misty1212):

oh except for the range

OpenStudy (anonymous):

I'm so bad at finding that.

OpenStudy (misty1212):

in this case it is real easy, since it has a vertical asympote and no horizontal asymptote that means it goes from \(-\infty\) to \(\infty\)

OpenStudy (misty1212):

you can write your answer as an interval \((-\infty, \infty)\) or just say "all real numbers"

OpenStudy (misty1212):

it is harder for some other ones, but not for this one

OpenStudy (anonymous):

Yay! Thanks so much:)))

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]